OpenStudy (anonymous):

Write an equation of the line tangent to the given circle at the given point. x^2+y^2=17; (1, 4)

OpenStudy (anonymous):

I understand how to graph and write the equations for circles, but I don't understand the question.

Directrix (directrix):

Recall from Geometry that a radius of the circle drawn to the point of tangency of a tangent to a circle is perpendicular to the tangent. So, that tells you something about the slope of the tangent line.

Directrix (directrix):

x^2+y^2=17 = ( x - 0)² + y - 0)² = (√17)² What is the slope of the line segment joining (0,0) and (1,4) ?

OpenStudy (anonymous):

4?

Directrix (directrix):